OpenStudy (anonymous):

Multiply (2x+5y)(2x-5y)

OpenStudy (harsimran_hs4):

\[(a + b)(c + d) = ac + ad + bc + bd\] can you solve now?

OpenStudy (anonymous):

u could use the rule (a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

(2x + 5y)(2x - 5y) use the foil method....(first, outside, inside, last) First : (2x)(2x) = 4x^2 Outside : (2x)(-5y) = -10xy Inside : (5y)(2x) = 10xy Last : (5y)(-5y) = - 25y^2 now we have ... 4x^2 - 10xy + 10xy + (-25y^2) 4x^2 - 25y^2 I think that is how to do this problem